OpenStudy (anonymous):

What is the answer? 15/16 + 2/8?

OpenStudy (anonymous):

15/16+4/16=19/16

OpenStudy (anonymous):

4/16 = 2*2/2*8

OpenStudy (anonymous):

denominator has to be the same

OpenStudy (anonymous):

So what is the real answer?

OpenStudy (anonymous):

19/16

OpenStudy (anonymous):

1.1875

OpenStudy (anonymous):

Lets first make the denominators equal 2/8 is equivalent to 4/16 (multiply the numerator and denominator by 2) So it is easy to add now 15/16 + 4/16 = (15+4)/16 = 19/16 This can also be expressed as 1 and 3/16